OpenStudy (wwwhhhaaattt?):

Find f(5) for this sequence: f(1) = 2 and f(2) = 5, f(n) = f(1) + f(2) + f(n - 1), for n > 2. f(5) = ______

OpenStudy (freckles):

first can you find f(3)?

OpenStudy (wwwhhhaaattt?):

@StudyGurl14

OpenStudy (wwwhhhaaattt?):

No, I suck at math

OpenStudy (freckles):

well f(n)=f(1)+f(2)+f(n-1) f(1) is 2 so I'm going to replace f(1) with 2 f(2) is 5 so I'm going to replace f(2) with 5 so we can write f(n) as f(n)=2+5+f(n-1) f(n)=7+f(n-1)<---we will just use this

OpenStudy (freckles):

To find f(3) you need to replace the n's with 3's

OpenStudy (freckles):

can you do that?

OpenStudy (freckles):